Check whether the following sequences (sn) are Cauchy, where (i)sn=1+2+3+...+n 4n3 + 3n (ii) sn 3n + n2 


1
Expert's answer
2021-03-08T18:34:49-0500

Solution:

We know that every cauchy sequence is convergent. So we check convergence for given sequences.

(i) "S_n=1+2+3+...+n=\dfrac{n(n+1)}2"

By series convergent test, we know that this series diverges.

Hence, this one is not cauchy sequence.

(ii) "S_n=\dfrac{4n^3+3n}{3n+n^2}"

"S_n=\dfrac{n^3(4+\frac{3}{n^2})}{n^2(\frac{3}{n}+1)}"

"S_n=\dfrac{n(4+\frac{3}{n^2})}{(\frac{3}{n}+1)}"

"\lim_{n\rightarrow \infty}S_n=\dfrac{\infty(4+0)}{(0+1)}=" diverges

Hence, this one also is not cauchy sequence.